Please can someone help me with these questions:
The sum of the roots of the equation x squared - 20x = 0 is_____________________.
The product of its roots is_________________.

Answer by stanbon(75887)

Comment:
The roots are:
(-b/(2a) + [sqrt(b^2-4ac]/(2a)
(-b/(2a) - [sqrt(b^2-4ac]/(2a)
----
If you add these you will get 2(-b/2a) = -b/a
=======================================================
If you multiply them them you will get:
(-b/2a)^2 - {[sqrt(b^2-4ac]/(2a)}^2
= (b^2/4a^2) - [(b^2-4ac)/(4a^2)]
---
= +4ac/(4a^2)
= c/a
=======================================================
The sum of the roots of the equation x squared - 20x = 0 is_____________________.
The sum of the roots of a quadratic are 2(-b/(2a) = -b/a
Your Problem:
a = 1 ; b = -20 ; c = 0
Sum of roots = -(-20/1) = 20
-------------------------------------------
The product of its roots is_________________.
The product of the roots is c/a = 0/1 = 0

================================================
Cheers,
Stan H.